Find a polynomial that represents the area of a circle whose radius is b+1 meters. Using the value of pi=3.14, what would be the radius if the area was to equate 4pi square meters?
Answer by thechamp1011(19)

WE KNOW THE FORMULA FOR AREA OF THE CIRCLE IS: A=Pi (r)^2
if r =b+1......................................A=pi(b+1)^2
A=pi (b^2+2b+1)

FOR THE SECOND PART OF THE PROBLEM: A= PI (R)^2
SINCE A=4PI.........................4PI=PI (R)^2
DIVIDE BOTH SIDE BY PI..............4PI/PI=PI/PI (R)^2
....................................4=(r)^2
TO FIND R, TAKE SQRT ON BOTH SIDE....r=+-2